OpenStudy (anonymous):

is 2x-5y=10 and -2x=3y+6 infinite or no solution

OpenStudy (anonymous):

i got no solution

OpenStudy (anonymous):

Is it a system of equations? If so, simple to solve x = -15/8 and y = -3/4

OpenStudy (anonymous):

yup

OpenStudy (atlas):

y=-2 and x=0 The system has a finite solution

OpenStudy (atlas):

Add the corresponding left hand and right hand sides of both the equations

OpenStudy (anonymous):

I agree with atlas. It has a finite solution (y = -2, x=0). Here's a detailed solution: You have these 2 equations: \[Eq 1: 2x - 5y = 10\]\[Eq 2: -2x = 3y + 6\] Now, subtract 3y to both sides of Eq2 to get: \[Eq 1: 2x - 5y = 10\]\[Eq 2: -2x - 3y = 6\] Add the corresponding sides of the equations: \[(2x + (-2x)) + ((-5y) + (-3y)) = 10 + 6\] and you'll get: \[-8y = 16 \rightarrow y = -2\] To get x, just substitute -2 to y in one of the equations, say in Eq1: \[2x -5(-2) = 10\]\[2x + 10 = 10\]\[2x = 0\]\[x =0\] You'll get the same result even if you do the substitution in Eq2. :)

OpenStudy (anonymous):

kcpass ,u have given answer with accurate precision.